OpenStudy (anonymous):

Find the equation, in standard form, of the line passing through the points (5,1) and (-2,3).

OpenStudy (anonymous):

use \[\frac{ y-y _{1} }{ x-x _{1} }=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (anonymous):

once you plug in the numbers, what do you do?

OpenStudy (anonymous):

try to simplify and get y=mx+c where m=slope c=y-intercept

OpenStudy (triciaal):

you are given 2 points on the line use this to find the slope the slope is the change in y / corresponding change in x m = (y2 -y1) / (x2 -x1) once you have the slope use one of the points to find the y-intercept, b now you have m and b you can write the general equation. for this problem you have (5,1) and (-2, 3) the slope ,m = 3-1/-2-5 = 2/-7 using (3, -2) 3 = -2/7(-2) + b b = 3-4/7 =(21-4)/7 y = -2/7(x) + 17/7 multiply everything by 7 and rewrite to the form ax + by = c

OpenStudy (triciaal):

2x + 7y = 17